OpenStudy (anonymous):

A building measures 50 feet in length and w feet in width.Which of the folowing is an expression for the perimeter of the building a.50/w b.50w c.50+w d.100+2w

OpenStudy (anonymous):

perimeter = 2(L+B) 2(50 + W) = 100 + 2W